# 6G: The Intelligent Network of Everything – A Comprehensive Vision, Survey, and Tutorial

### v2

@@ VIII-F3 Wireless Energy Transfer
    In the near-field category, energy is transferred by inductive coupling via magnetic fields or capacitive coupling via electric fields [539, 538]. In the near-field technologies, the distances over which the energy can be efficiently transferred are short. Typical commercial applications of the near-field WET include wireless charging of mobile phones, tablets, smart watches, electric toothbrushes, electronic medical implants, and electric vehicles, to mention a few.
@@
    Over the years, WET with microwave energy beamforming has been extensively studied. Traditional research has mainly focused on the RF aspects, while recent studies have expanded to signal and system designs as well, with a special focus on integrating WET into future wireless networks [536]. In this respect, the concept of SWIPT started to gain interest in the 2010s [544]. Currently, WET is being studied for 6G networks, with a special emphasis on energizing low-power IoT devices [537]. While there exist no large-scale commercial applications of the far-field WET, there are some in the near-field domain, such as wireless chargers for wireless/mobile devices, electronic medical implants, and electric vehicles. Although WET has been extensively studied over the decades, with some commercialized products, it is far from maturity. From a regulatory perspective, ITU-R published its first report on radio-frequency WET in 2016, with an updated version in 2021 [543]. This report focuses on the possible applications of WET via radio waves, corresponding technologies, and candidate spectrum bands.
@@
  • Literature and Future Directions: WET has been studied for decades in the literature. Traditional research has mainly focused on the RF design aspects, with a special emphasis on the energy receiver [545]. Since the early 2010s, signal and system designs have gained more interest [545], particularly in the context of integrating WET into future wireless networks [536]. In the literature, there are lots of survey articles which cover a wide range of aspects on WET [546, 544, 545, 538, 537, 536, 547, 548, 549, 550, 551, 552, 531, 553]. An overview of wireless powered communication networks was presented in [546]. The focus was on basic architectures, key technologies, and future extensions. The concept of SWIPT was meticulously reviewed in [544]. The studied aspects include basics, interference exploitation, emerging technologies, and future challenges. The work [545] reviewed WIPT from the perspectives of RF, signal, and system designs. In [538], a survey was provided on non-radiative near-field WET, covering fundamentals, challenges, metamaterials, and applications. In [537], the authors studied massive WET for sustainable IoT in the 6G era. A detailed discussion was given on the system architecture, applications, technological enablers, energy beamforming schemes, and future directions.
@@
    TABLE X: Summary of energy-aware technologies for 6G
@@
VIII-G Energy-Aware Technologies for 6G
As the absolute energy consumption is constantly growing, it is of utmost importance to further improve the energy efficiency of mobile networks. In this respect, developing more efficient energy management technologies is one of the main goals and challenges of 6G. In this section, we focus on three energy-aware technologies: green networks, energy harvesting (EH), and backscatter communications. These technologies are summarized in Table X.
@@ VIII-G1 Green Networks
  • Literature and Future Directions: Recently, several surveys have been published on green communication and networking for 6G, reviewing the latest research progress and related literature [562, 555, 563, 554, 557, 564, 565, 558]. In [562], a survey was provided on green 6G networks. The focus is on network architectures (space-air-ground-sea, intelligence network, new network protocol stack) and promising technologies (THz/VLC, EH, molecular/quantum communication, blockchain-based security, intelligent materials). In [555], the authors discussed greener PHY layer technologies for 6G, analyzing the joint energy-spectral efficiency design in NOMA and waveform overlapping multiple access frameworks. The work [563] reviewed green-UAV communications for 6G in terms of power consumption models, trends, enabling techniques, applications, and open research problems.
@@ VIII-G2 Energy Harvesting
    Since RF-EH is still far from its potential, significant research efforts are needed in the future. Therefore, it is vital to address the main challenges mentioned earlier. For instance, improvements in the conversion efficiency and energy consumption of EH devices call for advanced antenna and circuit designs. Further studies are required to obtain and maintain an appropriate level of harvested power density in the devices, especially in mobile environments. Moreover, the locations of dedicated energy transmitters need to be properly determined to ensure fairness among EH devices. More experiments are also required in realistic environments to validate the performance of RF-EH systems. Further details of the aforementioned issues and future directions can be found in [539, 566].
@@ VIII-G3 Backscatter Communications
    The ambient backscatter systems exploit existing wireless systems, such as cellular networks, WiFi, and television broadcasts as the RF sources [542]. In contrast to the bistatic approach, ambient systems do not require any dedicated spectrum or RF sources for communication, leading to improved spectral, power, and cost efficiency. However, ambient RF signals are uncontrollable and unpredictable, leading to more complicated system design and unstable performance. Nevertheless, the ambient backscatter communication concept is the most evolved and promising one to be utilized in 6G for low-power, low-complexity, and low-cost communication scenarios, enabling applications such as energy-constrained battery-free IoT networks [577, 578]. A comprehensive overview of the backscatter communications systems is presented in [542].
@@
    TABLE XI: Summary of end-device-oriented technologies for 6G
@@
  • Literature and Future Directions: Modern backscatter research began to gain notable attention after the introduction of the bistatic [579] and ambient [580] backscatter concepts in the early 2010s. Since then, backscatter communication has been studied from various perspectives in the literature. This research can be divided into three main categories: fundamentals, emerging systems, and applications. The main topics in fundamental research include channel coding, modulation, channel modeling, channel estimation, detection, resource allocation, multiple access, performance analysis, and PHY layer security. Literature reviews of these topics can be found in [542, 586, 585, 578].
@@
VIII-H End-Device-Oriented Communication Technologies for 6G
6G is expected to support massive amounts of wireless devices, vehicles, and drones, thereby setting diverse requirements for mobile networks. In this picture, D2D, V2X, and cellular-connected UAV communications will play key roles. Each of these end-device-related technologies is discussed below and summarized in Table XI.
@@ VIII-H1 D2D Communications
  • Literature and Future Directions: In the literature, the early research of D2D communications dates back to the late 2000s [595]. The special focus was first on 4G LTE use cases [593] but it shifted to 5G [596] around the mid-2010s and has recently turned toward 6G [591]. Over the years, a broad range of D2D topics has been examined, such as device discovery, mode selection, interference management, power control, resource allocation, relaying, mobility management, and security/privacy [590, 597]. Recent survey articles have reviewed the latest literature, state-of-the-art solutions, recent advances, and open challenges of D2D communications [597, 598, 591, 599, 600, 601, 602, 603]. These surveys discussed D2D in terms of 5G [597], 6G [598, 591, 599], cognitive D2D [600], resource management [601], state-of-the-art solutions [602], and AI-based resource allocation [603].
@@
VIII-H2 V2X Communications
@@
    In the standardization-wise, there are two types of V2X communication technologies, i.e., IEEE’s WiFi-based V2X, known as Dedicated Short-Range Communication (DSRC), and 3GPP’s cellular-based V2X (C-V2X) [155]. WiFi-based DSRC, which supports V2V and V2I technologies, does not require any communication network infrastructure to form ad hoc-type connections among vehicles and between vehicles and traffic infrastructure. The shortcoming of WiFi-based V2X communications is that strict service requirements cannot be guaranteed due to unlicensed spectrum operation. In addition to V2V and V2I, C-V2X supports V2N connectivity, which is particularly important for enabling future intelligent vehicular and transportation systems.
@@
  • Opportunities and Challenges: The potential of next-generation V2X communications is to advance the connectivity, intelligence, autonomy, and automation of large-scale vehicular systems and individual vehicles; traffic tracking, monitoring, management, cooperation, and efficiency; cooperation between different objects and entities in the framework of vehicular systems; and ubiquitous road safety. From a broader perspective, V2X communication is seen as a key technology to enable intelligent vehicular and transportation systems in the 6G era. To realize this vision in practice, V2X must rely on a wide range of emerging technologies. Some potential enablers include a flexible utilization of a broad spectrum range (from sub-6 GHz to THz frequencies), pervasive use of AI/ML (from system architecture to air interface), coordinated utilization of 3D network architecture (ground-air-space), and broad usage of beyond-communication technologies (computing, sensing, localization, and control) [155]. While these technologies offer novel opportunities for V2X communications, they also introduce diverse challenges for the corresponding network design. The survey [155] provides a detailed discussion of the key enabling technologies for 6G-V2X, discussing their opportunities and challenges.
@@
VIII-H3 Cellular-Connected UAV Communications
@@
    From a communication perspective, the new concept of UAV-to-everything (U2X) communications, introduced in [619], is at the center of future cellular UAV connectivity. A generic definition of U2X communications is the connectivity between UAVs and other nodes in a cellular network, such as BSs, ground-devices, and other UAVs [619]. The main connectivity modes of U2X include UAV-to-network, UAV-to-UAV, and UAV-to-device communications [619]. Other potential modes are UAV-to-satellite and UAV-to-vehicle communications. Ultimately, U2X communications aim to take advantage of all layers of the network, from ground to space and from cellular to vehicular. In particular, U2X communication is seen as an enabler for IoU [619].
@@
    In Release 15, the work on cellular-operated UAVs started with a study item on the required enhancements of LTE networks to support UAVs, resulting in TR 36.777 [620]. In Release 16, the work continued with TR 22.829, which identified UAV use cases and the required enhancements for 3GPP networks, calling for 5G support [620]. Moreover, TS 22.125 defined the requirements for UAS operation over 3GPP networks [620]. In Release 17, TR 23.754 and TR 23.755 studied mechanisms to support connectivity, identification, and tracking in UAS, as well as the architectural requirements and solutions to support UAS applications, respectively [620]. Release 18 provided further enhancements for the NR support of UAVs, especially in terms of identification, reporting, broadcasting, and beamforming [162]. A comprehensive survey on the 3GPP standardization efforts for UAS is given in [620].
@@
    TABLE XII: Summary of service-oriented technologies for 6G
@@
  • Literature and Future Directions: The research on cellular-connected UAVs began gaining popularity in the 2010s. Diverse aspects and issues have been addressed and solutions proposed. In the past few years, researchers have conducted numerous surveys, reviewing topics that range from fundamentals to standardization, recent advancements, open challenges, and future research guidelines [622, 623, 619, 624, 620, 626, 625, 162, 627, 628, 629, 164, 165, 166]. In [622], the article reviewed cellular-connected UAV operations in terms of communication/spectrum requirements, design aspects, potential technologies, and future directions. The authors in [623] presented a survey on cellular UAV communications, with the focus on UAV types, standardization, UAV-BSs, prototyping/field testing, regulations, security, and future research areas. The work [619] introduced a novel concept of U2X communications from the perspectives of fundamentals, key techniques, reinforcement learning framework, and future extensions. Cellular-operated UAVs were explored in [624], covering applications/use cases, challenges, 5G/B5G innovations, trials/prototyping, standardization, and future research guidelines.
    The paper [620] reviewed 3GPP standardization of cellular UAS support, and introduced the key research drivers. In [626], the authors discussed enabling technologies, standardization efforts, security, and open issues for advanced cellular UAV operations. The authors envisioned an evolution road for cellular UAV connectivity from 5G to 6G in [625]. Many aspects were discussed, including 5G NR and beyond, sub-6 GHz massive MIMO, mmWave/THz, AI/ML, NTNs, RISs, UAV-to-UAV communications, and future challenges. The study [627] discussed cellular UAV support using THz frequencies. An overview of UAV clients for beyond 5G was presented in [162]. The paper [628] reviewed advanced air mobility in cellular networks, with a particular focus on use cases and beamforming techniques. In [629], the authors explored the use of unmanned vehicles in 6G networks. The surveys [164, 165, 166] explored cellular-supported cooperative UAV swarms.
@@
VIII-I Service-Oriented Technologies for 6G
It is anticipated that 6G will significantly expand the services of mobile networks, especially in vertical domains. Private networks will be at the core of this evolution, offering customized services for vertical clients and creating novel business opportunities. Private networks are reviewed in the following and summarized in Table XII.
@@ VIII-I1 Private Networks
  • Past and Present: Traditionally, mobile networks have been mostly public, offering services to conventional mobile subscribers. Private networks were first introduced in 4G. However, the usage has been rather marginal, partly due to the limitations of 4G. 5G private networks, also known as non-public networks in 3GPP, were introduced in Release 16 [630]. Compared to 4G, 5G provides significantly better performance and capabilities for private networks, opening up new service opportunities, especially in industrial environments [631]. In the 5G NR standards, private networks are supported in two categories, i.e., public network integrated and stand-alone non-public networks [630]. The former relies on a PLMN and the latter performs independently. Full control and independence to customize the stand-alone network according to the customer’s needs comes at the price of higher capital and operational costs [630]. Further details on the standardization of 5G non-public networks can be found in [630]. In the big picture, 5G is paving the way for a paradigm shift where mobile networks are not only used for serving public subscribers but also providing versatile services for a wide range of vertical industries.
@@
    Although private networks have been included in the 5G standardization and commercial solutions exist around the world, they are still at a rather early evolutionary stage and far from their true potential. There are diverse technological obstacles on the way toward freeing the full potential of private networks. First, private networks must fulfill the requirements of customer verticals, such as tailored services, guaranteed QoS, accurate coverage, protected data, and secure networking [630]. Additionally, there is a constant need to further enhance the performance and capabilities of private networks. Whereas 5G evolution can provide rather limited enhancements, 6G is expected to take private networks to the next level, with extreme performance, pervasive AI/ML, and integrated beyond-communication technologies. In addition to opportunities, 6G will introduce new challenges as well. In this picture, the main issues are related to resource management, spectrum use, AI/ML-assisted network optimization, integrated computation, control, and sensing, fronthaul/backhaul, data privacy, and network security. In [630, 632, 633, 634], recent surveys provide detailed discussions on the main challenges of private networks.
    TABLE XIII: Summary of security technologies for 6G
@@
    The paper [631] discussed private 5G networks in an industrial context. The covered topics include industrial networking demands, 5G opportunities for industrial wireless, functional architecture of 5G private networks, industrial usage scenarios, licensed/unlicensed spectrum operation, network design challenges, and standardization efforts. In [635], 5G and beyond private networks were explored in terms of application scenarios, standardization aspects, operator models, and technological enablers. The authors of [630] reviewed 5G non-public networks from a standardization perspective, focusing on enterprise customer requirements, enabling technology solutions, 3GPP standardization, single- and multi-site network scenarios, and implementation challenges.
@@
VIII-J Security Technologies for 6G
Since future society will be highly dependent on 6G, it is of utmost importance to make 6G networks trustworthy and secure. A special attention needs to be paid to privacy as well, as data is becoming more abundant. Therefore, a comprehensive network security architecture is required, as discussed below and summarized in Table XIII.

### v3

@@ IX-F3 WIRELESS ENERGY TRANSFER
    In the near-field category, energy is transferred by inductive coupling via magnetic fields or capacitive coupling via electric fields [548]. In the near-field technologies, the distances over which the energy can be efficiently transferred are short. Typical commercial applications of the near-field WET include wireless charging of mobile phones, tablets, smart watches, electric toothbrushes, electronic medical implants, and electric vehicles, to mention a few.
@@
    Over the years, WET with microwave energy beamforming has been extensively studied. Traditional research has mainly focused on the RF aspects, while recent studies have expanded to signal and system designs as well, with a special focus on integrating WET into future wireless networks [546]. In this respect, the concept of SWIPT started to gain interest in the 2010s [553]. Currently, WET is being studied for 6G networks, with a special emphasis on energizing low-power IoT devices [547]. While there exist no large-scale commercial applications of the far-field WET, there are some in the near-field domain, such as wireless chargers for wireless/mobile devices, electronic medical implants, and electric vehicles. Although WET has been extensively studied over the decades, with some commercialized products, it is far from maturity. From a regulatory perspective, ITU-R published its first report on radio-frequency WET in 2016, with an updated version in 2021 [552]. This report focused on the possible applications of WET via radio waves, corresponding technologies, and candidate spectrum bands.
@@
  • Literature and Future Directions: In the literature, there are lots of survey articles which have reviewed a wide range of WET aspects [554, 548, 547, 546, 555, 556, 557, 558, 559, 560, 541, 561]. The work [554] reviewed WIPT from the perspectives of RF, signal, and system designs. In [548], a survey was provided on non-radiative near-field WET, covering fundamentals, challenges, metamaterials, and applications. In [547], the authors studied massive WET for sustainable IoT in the 6G era. A detailed discussion was given on the system architecture, applications, technological enablers, energy beamforming schemes, and future directions.
@@
    TABLE XII: Summary of energy-aware technologies for 6G
@@
IX-G ENERGY-AWARE TECHNOLOGIES FOR 6G
As the absolute energy consumption is constantly growing, it is of utmost importance to further improve the energy efficiency of mobile networks. In this respect, developing more efficient energy management technologies is one of the main goals and challenges of 6G. In this section, we focus on three energy-aware technologies: green networks, energy harvesting (EH), and backscatter communications. These technologies are summarized in Table XII.
@@ IX-G1 GREEN NETWORKS
  • Literature and Future Directions: Recently, several surveys have been published on green communication and networking for 6G, reviewing the latest research progress and related literature [570, 563, 571, 562, 565, 572, 573, 566]. In [570], a survey was provided on green 6G networks. The focus was on network architectures (space-air-ground-sea, intelligence network, new network protocol stack) and promising technologies (THz/VLC, EH, molecular/quantum communication, blockchain-based security, intelligent materials). In [563], the authors discussed greener PHY layer technologies for 6G, analyzing the joint energy-spectral efficiency design in NOMA and waveform overlapping multiple access frameworks. The work [571] reviewed green-UAV communications for 6G in terms of power consumption models, trends, enabling techniques, applications, and open research problems.
@@ IX-G2 ENERGY HARVESTING
    Since RF-EH is still far from its potential, significant research efforts are needed in the future. Therefore, it is vital to address the main challenges mentioned earlier. For instance, improvements in the conversion efficiency and energy consumption of EH devices call for advanced antenna and circuit designs. Further studies are required to obtain and maintain an appropriate level of harvested power density in the devices, especially in mobile environments. Moreover, the locations of dedicated energy transmitters need to be properly determined to ensure fairness among EH devices. More experiments are also required in realistic environments to validate the performance of RF-EH systems. Further details of the aforementioned issues and future directions can be found in [575].
IX-G3 BACKSCATTER COMMUNICATIONS
@@
    The ambient backscatter systems exploit existing wireless systems, such as cellular networks, WiFi, and television broadcasts as the RF sources [551]. In contrast to the bistatic approach, ambient systems do not require any dedicated spectrum or RF sources for communication, leading to improved spectral, power, and cost efficiency. However, ambient RF signals are uncontrollable and unpredictable, leading to more complicated system design and unstable performance. Nevertheless, the ambient backscatter communication concept is the most evolved and promising one to be utilized in 6G for low-power, low-complexity, and low-cost communication scenarios, enabling applications such as energy-constrained battery-free IoT networks [586, 587]. A comprehensive overview of the backscatter communications systems was presented in [551].
@@
    TABLE XIII: Summary of end-device-oriented technologies for 6G
@@
  • Literature and Future Directions: Modern backscatter research began to gain notable attention after the introduction of the bistatic and ambient backscatter concepts in the early 2010s. Since then, backscatter communication has been studied from various perspectives in the literature. This research can be divided into three main categories: fundamentals, emerging systems, and applications. The main topics in fundamental research included channel coding, modulation, channel modeling, channel estimation, detection, resource allocation, multiple access, performance analysis, and PHY layer security. Literature reviews of these topics can be found in [551, 595, 594, 587].
@@
IX-H END-DEVICE-ORIENTED COMMUNICATION TECHNOLOGIES FOR 6G
6G is expected to support massive amounts of wireless devices, vehicles, and drones, thereby setting diverse requirements for mobile networks. In this picture, D2D, V2X, and cellular-connected UAV communications will play key roles. Each of these end-device-related technologies is discussed below and summarized in Table XIII.
@@ IX-H1 D2D COMMUNICATIONS
  • Literature and Future Directions: In the literature, the early research focus of D2D communications was first on 4G LTE use cases [604, 602] but it shifted to 5G [605] around the mid-2010s and has recently turned toward 6G [600]. Over the years, a broad range of D2D topics has been examined, such as device discovery, mode selection, interference management, power control, resource allocation, relaying, mobility management, and security/privacy [599, 606]. Recent survey articles have reviewed the latest literature, state-of-the-art solutions, recent advances, and open challenges of D2D communications [606, 607, 600, 608, 609, 610, 611, 612]. These surveys discussed D2D in terms of 5G [606], 6G [607, 600, 608], cognitive D2D [609], resource management [610], state-of-the-art solutions [611], and AI-based resource allocation [612].
@@
Figure 45: Four main types of V2X communications.
IX-H2 V2X COMMUNICATIONS
@@
    In the standardization-wise, there are two types of V2X communication technologies, i.e., IEEE WiFi-based V2X, known as Dedicated Short-Range Communication (DSRC), and 3GPP cellular-based V2X (C-V2X) [173]. WiFi-based DSRC, which supports V2V and V2I technologies, does not require any communication network infrastructure to form ad hoc-type connections among vehicles and between vehicles and traffic infrastructure. The shortcoming of WiFi-based V2X communications is that strict service requirements cannot be guaranteed due to unlicensed spectrum operation. In addition to V2V and V2I, C-V2X supports V2N connectivity, which is particularly important for enabling future intelligent vehicular and transportation systems.
@@
  • Opportunities and Challenges: The potential of next-generation V2X communications is to advance the connectivity, intelligence, autonomy, and automation of large-scale vehicular systems and individual vehicles; traffic tracking, monitoring, management, cooperation, and efficiency; cooperation between different objects and entities in the framework of vehicular systems; and ubiquitous road safety. From a broader perspective, V2X communication is seen as a key technology to enable intelligent vehicular and transportation systems in the 6G era. To realize this vision in practice, V2X must rely on a wide range of emerging technologies. Some potential enablers include a flexible utilization of a broad spectrum range (from sub-6 GHz to THz frequencies), pervasive use of AI/ML (from system architecture to air interface), coordinated utilization of 3D network architecture (ground-air-space), and broad usage of beyond-communication technologies (computing, sensing, localization, and control) [173]. While these technologies offer novel opportunities for V2X communications, they also introduce diverse challenges for the corresponding network design. The survey [173] provided a detailed discussion of the key enabling technologies for 6G-V2X, discussing their opportunities and challenges.
@@
IX-H3 CELLULAR-CONNECTED UAV COMMUNICATIONS
@@
    From a communication perspective, the new concept of UAV-to-everything (U2X) communications, introduced in [628], is at the center of future cellular UAV connectivity. A generic definition of U2X communications is the connectivity between UAVs and other nodes in a cellular network, such as BSs, ground-devices, and other UAVs [628]. The main connectivity modes of U2X include UAV-to-network (U2N), UAV-to-UAV (U2U), and UAV-to-device (U2D) communications [628]. Other potential modes are UAV-to-satellite (U2S) and UAV-to-vehicle (U2V) communications. Ultimately, U2X communications aim to take advantage of all layers of the network, from ground to space and from cellular to vehicular. In particular, U2X communication is seen as an enabler for IoU [628]. A schematic illustration of potential U2X types is shown in Figure 47.
    Figure 47: Five potential types of U2X communications.
@@
    In Release 15, the work on cellular-operated UAVs started with a study item on the required enhancements of LTE networks to support UAVs, resulting in TR 36.777 [629]. In Release 16, the work continued with TR 22.829, which identified UAV use cases and the required enhancements for 3GPP networks, calling for 5G support [629]. Moreover, TS 22.125 defined the requirements for UAS operation over 3GPP networks [629]. In Release 17, TR 23.754 and TR 23.755 studied mechanisms to support connectivity, identification, and tracking in UAS, as well as the architectural requirements and solutions to support UAS applications, respectively [629]. Release 18 provided further enhancements for the NR support of UAVs, especially in terms of identification, reporting, broadcasting, and beamforming [180]. A comprehensive survey on the 3GPP standardization efforts for UAS was given in [629].
@@
    TABLE XIV: Summary of service-oriented technologies for 6G
@@
  • Literature and Future Directions: In recent years, researchers have conducted numerous surveys on cellular-connected UAVs, reviewing topics that range from fundamentals to standardization, recent advancements, open challenges, and future research guidelines [631, 632, 628, 633, 629, 635, 634, 180, 636, 637, 638, 182, 183, 184]. In [631], the article reviewed cellular-connected UAV operations in terms of communication/spectrum requirements, design aspects, potential technologies, and future directions. The authors in [632] presented a survey on cellular UAV communications, with the focus on UAV types, standardization, UAV-BSs, prototyping/field testing, regulations, security, and future research areas. The work [628] introduced a novel concept of U2X communications from the perspectives of fundamentals, key techniques, reinforcement learning framework, and future extensions. Cellular-operated UAVs were explored in [633], covering applications/use cases, challenges, 5G/B5G innovations, trials/prototyping, standardization, and future research guidelines.
    The paper [629] reviewed 3GPP standardization of cellular UAS support, and introduced the key research drivers. In [635], the authors discussed enabling technologies, standardization efforts, security, and open issues for advanced cellular UAV operations. The authors envisioned an evolution road for cellular UAV connectivity from 5G to 6G in [634]. Many aspects were discussed, including 5G NR and beyond, sub-6 GHz massive MIMO, mmWave/THz, AI/ML, NTNs, RISs, UAV-to-UAV communications, and future challenges. The study [636] discussed cellular UAV support using THz frequencies. An overview of UAV clients for beyond 5G was presented in [180]. The paper [637] reviewed advanced air mobility in cellular networks, with a particular focus on use cases and beamforming techniques. In [638], the authors explored the use of unmanned vehicles in 6G networks. The surveys [182, 183, 184] examined cellular-supported cooperative UAV swarms.
@@
IX-I SERVICE-ORIENTED TECHNOLOGIES FOR 6G
It is anticipated that 6G will significantly expand the services of mobile networks, especially in vertical domains. Private networks will be at the core of this evolution, offering customized services for vertical clients and creating novel business opportunities. Private networks are reviewed in the following and summarized in Table XIV.
@@ IX-I1 PRIVATE NETWORKS
  • Past and Present: Traditionally, mobile networks have been mostly public, offering services to conventional mobile subscribers. Private networks were first introduced in 4G. However, the usage has been rather marginal, partly due to the limitations of 4G. 5G private networks, also known as non-public networks in 3GPP, were introduced in Release 16 [639]. Compared to 4G, 5G provides significantly better performance and capabilities for private networks, opening up new service opportunities, especially in industrial environments [640]. In the 5G NR standards, private networks are supported in two categories, i.e., public network integrated and stand-alone non-public networks [639]. The former relies on a PLMN and the latter performs independently. Full control and independence to customize the stand-alone network according to the needs of the customer comes at the price of higher capital and operational costs [639]. Further details on the standardization of 5G non-public networks can be found in [639]. In the big picture, 5G is paving the way for a paradigm shift where mobile networks are not only used for serving public subscribers but also providing versatile services for a wide range of vertical industries.
@@
    Although private networks have been included in the 5G standardization and commercial solutions exist around the world, they are still at a rather early evolutionary stage and far from their true potential. There are diverse technological obstacles on the way toward freeing the full potential of private networks. First, private networks must fulfill the requirements of customer verticals, such as tailored services, guaranteed QoS, accurate coverage, protected data, and secure networking [639]. Additionally, there is a constant need to further enhance the performance and capabilities of private networks. Whereas 5G evolution can provide rather limited enhancements, 6G is expected to take private networks to the next level, with extreme performance, pervasive AI/ML, and integrated beyond-communication technologies. In addition to opportunities, 6G will introduce new challenges as well. In this picture, the main issues are related to resource management, spectrum use, AI/ML-assisted network optimization, integrated computation, control, and sensing, fronthaul/backhaul, data privacy, and network security. In [639, 641, 642, 643], recent surveys provided detailed discussions on the main challenges of private networks.
    TABLE XV: Summary of security technologies for 6G
@@
    The paper [640] discussed private 5G networks in an industrial context. The main topics included industrial networking demands, 5G opportunities for industrial wireless, functional architecture of 5G private networks, industrial usage scenarios, licensed/unlicensed spectrum operation, network design challenges, and standardization efforts. In [644], 5G and beyond private networks were explored in terms of application scenarios, standardization aspects, operator models, and technological enablers. The authors of [639] reviewed 5G non-public networks from a standardization perspective, focusing on enterprise customer requirements, enabling technology solutions, 3GPP standardization, single- and multi-site network scenarios, and implementation challenges.
@@
IX-J SECURITY TECHNOLOGIES FOR 6G
Since future society will be highly dependent on 6G, it is of utmost importance to make 6G networks trustworthy and secure. A special attention needs to be paid to privacy as well, as data is becoming more abundant. Therefore, a comprehensive network security architecture is required, as discussed below and summarized in Table XV.